# Inverse Constitutional AI: Compressing Preferences into Principles

### v1

@@ Appendix A Additional experimental results
Appendix B Prompts
@@
Note that prompts are generally separated into two messages, a system message and a user message. We use the following format for all prompts (based on AlpacaEval’s formatting): <|im_start|> and <|im_end|> denote the start and end of a message, followed by the message type (system or user) and the content.
B.1 Principle generation
@@ B.3 Constitution evaluation
B.4 Non-constitutional baseline
We also evaluate the preferences the language model expresses when not given a constitution to follow, i.e., the biases inherent in the trained model when asked to select the “best” output. We adapted two of the default prompts from AlpacaEval for this purpose by removing references to an “instruction”, as this is not present in all pairwise ranking datasets. We selected the alpacaeval_gpt4_turbo_fn and chatgpt_fn prompts as they were reported to have the highest human agreement rate for the gpt-4-turbo and gpt-3.5-turbo models, respectively, while also being below an (estimated) price of 6\$/1k examples.  [^10] https://github.com/tatsu-lab/alpaca_eval/tree/v0.6.2/src/alpaca_eval/evaluators_configs
@@ B.4 Non-constitutional baseline
Appendix C Constitutions
The following lists examples of full constitutions generated for each dataset and model combination in our experiments. To provide an unbiased view of the generated constitutions, we show the constitution with the highest, median, and lowest performance on the test set. Since even numbers of seeds are used, we chose the worse-performing constitution as a tie-breaker for the median.
C.1 Synthetic datasets
@@
C.1.1 Aligned
Listing 9: Best constitution on the ‘aligned’ synthetic dataset.
@@ C.1.1 Aligned
Listing 10: Median constitution on the ‘aligned’ synthetic dataset.
@@
Listing 11: Worst constitution on the ‘aligned’ synthetic dataset.
@@
C.1.2 Orthogonal
Listing 12: Best constitution on the ‘orthogonal’ synthetic dataset.
@@ C.1.2 Orthogonal
Listing 13: Median constitution on the ‘orthogonal’ synthetic dataset.
@@
Listing 14: Worst constitution on the ‘orthogonal’ synthetic dataset.
@@
C.1.3 Unaligned
Listing 15: Best constitution on the ‘unaligned’ synthetic dataset.
@@ C.1.3 Unaligned
Listing 16: Median constitution on the ‘unaligned’ synthetic dataset.
@@
Listing 17: Worst constitution on the ‘unaligned’ synthetic dataset.
@@
C.2 AlpacaEval datasets
C.2.1 Aligned
Listing 18: Best constitution on the ‘aligned’ AlpacaEval dataset.
@@ C.2.1 Aligned
Listing 19: Median constitution on the ‘aligned’ AlpacaEval dataset.
@@
Listing 20: Worst constitution on the ‘aligned’ AlpacaEval dataset.
@@
C.2.2 Unaligned
Listing 21: Best constitution on the ‘unaligned’ AlpacaEval dataset.
@@ 4. Select the response that lists key takeaways clearly and concisely
Listing 22: Median constitution on the ‘unaligned’ AlpacaEval dataset.
@@
Listing 23: Worst constitution on the ‘unaligned’ AlpacaEval dataset.
@@ C.3 Chatbot Arena
C.3.1 User A
Listing 24: Best constitution on User A annotations.
@@ C.3.1 User A
Listing 25: Median constitution on User A annotations.
@@
Listing 26: Worst constitution on User A annotations.
@@
C.3.2 User B
Listing 27: Best constitution on User B annotations.
@@ C.3.2 User B
Listing 28: Median constitution on User B annotations.
@@
Listing 29: Worst constitution on User B annotations.
@@
Appendix D Numerical Results
In addition to the bar plots shown in the main paper, we provide a table of numerical results for all experiments here.
Dataset Model Annotator Mean Std Min Max
Orthogonal GPT-3.5 Turbo Constitutional 86.67% 8.43 73.33% 96.67%
Default 37.78% 2.72 33.33% 40.00%
Aligned GPT-3.5 Turbo Constitutional 92.22% 6.89 83.33% 100.00%
Default 100.00% 0.00 100.00% 100.00%
Unaligned GPT-3.5 Turbo Constitutional 84.44% 9.81 73.33% 100.00%
Default 0.00% 0.00 0.00% 0.00%
Table 2: Results for experiments on AlpacaEval data. Averaged over 6 random seeds.
Dataset Model Annotator Mean Std Min Max
Aligned GPT-3.5-Turbo Default 64.87% 1.16 63.08% 66.15%
Constitutional 67.44% 3.43 63.08% 72.31%
GPT-4o Default 72.56% 1.16 70.77% 73.85%
Constitutional 68.46% 3.19 63.08% 72.31%
Unaligned GPT-3.5-Turbo Default 35.90% 1.26 33.85% 36.92%
Constitutional 39.49% 3.32 35.38% 44.62%
GPT-4o Default 26.92% 1.61 24.62% 29.23%
Constitutional 66.41% 7.69 53.85% 72.31%
Dataset Model Annotator Mean Std Min Max
User B constitution 83.33% 5.27 75.00% 91.67%
User B constitution 94.44% 6.09 88.89% 100.00%
Dataset Model Annotator Mean Std Min Max
Unaligned GPT-4o Default 26.15% 1.26 24.62% 27.69%
Constitutional 70.00% 0.89 69.23% 70.77%
Claude-3-Opus Default 24.23% 0.77 23.08% 24.62%
Constitutional 58.85% 0.77 58.46% 60.00%
Claude-3-Haiku Default 36.92% 0.00 36.92% 36.92%
Constitutional 59.65% 0.00 59.65% 59.65%
Dataset Model Annotator Mean Std Min Max
Constitutional (n=3) 75.00% 14.10 63.33% 100.00%
Constitutional (n=5) 86.67% 8.43 73.33% 96.67%
Constitutional (n=10) 90.00% 10.11 70.00% 96.67%
Appendix E Cost estimates
In this section we estimate the cost of reproducing all experiments shown in this paper. All experiments were run using models via API access from OpenAI and Anthropic. Note that all estimates are subject to variability due to provider pricing as well as inherent variability of the length (and thus cost) of model outputs.
We estimate the remaining cost of experiments to be less than 5\$. Overall we thus estimate the total cost of re-running our experiments to approx. 52.30\$ in API costs. Note that the overall cost for running experiments in the context of this project was about 3 times this amount (approx. 156.90\$), due to failed runs and additional experimentation that did not fit into the scope of the paper.
Appendix F Synthetic data generation

### v2

@@ F.5 Hyperparameter sensitivity
F.6 Constitution transferability
We investigate the transferability of constitutions across different model families. Shown in Figure 9, the results indicate that the constitution generated by GPT-4o transfers well to models from the Claude family, Claude-3-Opus and Claude-3-Haiku.
Figure 9: Transferability of constitutions: results of transferring a GPT-4o generated constitution to other model family (Claude). We use the highest-performing unaligned constitution on the training set, from experiments shown in the unaligned plot in Figure 4. We test two additional models from the Claude model family, Claude-3-Opus and Claude-3-Haiku. Both are able to use GPT-4o’s generated constitution to reconstruct the test set annotations effectively, albeit to a lower standard than GPT-4o. Plots show mean and standard deviation using 4 seeds per annotator, all with the same constitution.
F.7 Results on large datasets
Figure 10 and Table 10 show the results of using the entire 648 samples in the cross-annotated AlpacaEval dataset in our experiments (AlpacaEval-Large, see Appendix A), instead of the 130 samples used in the original experiments. We use 324 samples for training and 324 for testing.
Figure 10: Scaling up experiments on the AlpacaEval unaligned dataset. We scale our original experiment up 5×5$\times$5 × to the entire 648 samples in the cross-annotated AlpacaEval dataset, instead of the 130 samples used in the original experiments. As before we split the dataset in half to obtain a test and training set, using 324 samples for training (generating the constitution) and 324 for testing. We also provide the original results for comparison.
F.8 Extended Baseline Discussion
We compare our method against several baselines, described in detail below. All results discussions are based on Tables 4, 5 and 10.
Default
These baseline annotators vary depending on the model used to run them (GPT-3.5-Turbo or GPT-4o) and are directly based on two annotator configurations leading in their model class in the AlpacaEval (AE) evaluator leaderboard, [^11] See https://github.com/tatsu-lab/alpaca_eval/tree/main/src/alpaca_eval/evaluators_configs chatgpt_fn (used with GPT-3.5-Turbo) and alpaca_eval_gpt4_turbo_fn (used with GPT-4o). We only make small tweaks to the prompts to fit our data format (described in Appendix Section G.4) and update the original GPT-4-Turbo model with the newer GPT-4o model for the latter configuration (as no GPT-4o-specific configuration was available). We made a careful trade-off between reported cost (less than 6\$/1000k annotations) and performance (best with their model, at their price points) for our baselines. In particular, alpaca_eval_gpt4_turbo_fn is reported to perform (68.1%) close to the top configuration (alpaca_eval_gpt4_fn, 71.0%) discussed above. It is not tailored to the datasets used in our experiments. Consequently, its performance is expected to be strong on datasets aligned with the model’s training data and weaker on unaligned datasets. To account for this, we include a flipped version of this baseline, where predicted preference labels are inverted.
Results. As expected, we see that the baselines perform strongly on datasets aligned with the base model’s learned preferences, but poorly on other datasets. This is an inherent limitation of such an annotator, as it has no ability to adapt to new data.
Default (flipped)
This variant of the Default baseline uses the same AlpacaEval prompts but flips the predicted preference labels. Note that such a manual adjustment works only in limited scenarios such as our unaligned datasets; the default annotator cannot generally adapt to dataset-specific characteristics.
Results. Similar to the Default annotator, this baseline performs well on a restricted selection of datasets — just the inverse of the Default annotator (the unaligned datasets as opposed to the aligned ones).
PopAlign
This baseline is adapted from the PopAlign method developed by Wang et al. (2024). We modify this method, originally created for data generation, to the pairwise preference annotation setting. Similar to our method, PopAlign generates principles to annotate response pairs. However, instead of generating a fixed constitution representing an entire dataset (as in our method), PopAlign dynamically generates principles for each response pair and then annotates the pair according to the same principles. The detailed prompt and how we adapt the method is included in Section G.5. Many of the principles PopAlign generates as part of this process are qualitatively similar to those found in ICAI constitutions, for example: “A good response should be accurate, relevant, and provide clear and practical information”. We make this PopAlign-based annotator available as an AlpacaEval annotator config as part of our public package. [^12] See https://github.com/rdnfn/icai/tree/main/data/annotator_configs.
Results. While this baseline, similar to ICAI, generates principles, these principles are generated on-the-fly and without access to training data with known annotator preferences. Hence, the principles generated by this baseline are always in-line with its own learned preferences and cannot adapt to a new dataset, resulting in performance comparable to the Default annotator (good on aligned, bad on unaligned datasets).
PairRM
This baseline uses the Pairwise Reward Model (PairRM) [^13] Available at https://huggingface.co/llm-blender/PairRM, our experiments use revision 5b880cc73776ac75a835b3e0bd5169bcb5be013b. by Jiang et al. (2023), a black-box pairwise preference model with 400 million parameters. It accepts a pair of output candidates and an instruction as input, jointly encoding them to produce scores that reflect relative quality. Unlike the other baselines and our method, PairRM provides deterministic scores rather than relying on language model sampling. Therefore, we report results for a single seed without standard deviation or extrema. Similar to the Default annotator, PairRM is not customized to the datasets in our experiments, potentially leading to weaker performance on unaligned datasets. To evaluate sample efficiency and fairness, we also include a tuned version of PairRM that is fine-tuned on the training data.
Results. Since this version of the reward model is not fine-tuned, it has no ability to adapt to a dataset (similar to the Default and Default (flipped) annotators). Its relative performance mirrors the Default annotator, therefore, performing well (on-par with the Default annotator) on aligned datasets and poorly on others. The reward model’s performance exceeds the Default annotator’s on the synthetic-orthogonal dataset, which is likely due to a chance preference on the data chosen to be orthogonal to the Default annotator’s preferences.
PairRM (tuned)
This baseline uses the PairRM model fine-tuned on the training data prior to testing. Fine-tuning is performed for up to five additional epochs and a batch size of 1, with validation accuracy used to select the best model. The training data matches the data used to generate the constitution in our method, with a fraction of the training data (10 for AlpacaEval, 32 for AlpacaEval Large) reserved for validation. For synthetic data experiments, the model is tested on the same data used for fine-tuning, as separate test or validation sets are unavailable for this small dataset. This leads to overfitting and affects generalizability, which is less critical for our method, where interpretability is the primary focus and quantitative results are secondary. For fairness, the same (non-split) procedure is applied to PairRM. However, the reported performance on synthetic datasets likely overestimates the model’s capability on unseen data.
Results. PairRM is the only baseline that can, like ICAI, use training data to adapt to a new dataset. This is reflected in its reconstruction ability, generally exceeding the one of the non-fine-tuned version, especially on unaligned and orthogonal datasets. We observe that this ability to adapt is limited, however, as reflected in the model’s sub-par performance on the AlpacaEval unaligned setting. This is likely due to the model’s sensitivity to hyperparameters as well as the limited training data, which is completely opposed to the model’s (much larger) pretraining data.
F.9 Numerical Results
We provide full numerical results in table format for our experiments. Tables 4 and 5 show the numerical results for the core experiments on the synthetic and AlpacaEval datasets, respectively, featuring an extended set of baselines. Further, Tables 6 and 7 show the numerical results for the personalized experiments on the Chatbot Arena and PRISM datasets, Table 8 shows the results for the cross-model experiments, and Table 9 shows the results for the hyperparameter sensitivity experiments. Finally, Table 10 shows the results for the scaling experiments on AlpacaEval data. Experiments that were already discussed using a table previously are not repeated here.
Dataset Model Annotator Mean Std Min Max
Orthogonal GPT-3.5 Turbo Constitutional 86.67% 8.43 73.33% 96.67%
Default 37.78% 2.72 33.33% 40.00%
Default (flipped) 62.22% 1.72 60.00% 63.33%
PopAlign 38.89% 1.72 36.67% 40.00%
PairRM 73.33%
PairRM (tuned) 100.00%
Aligned GPT-3.5 Turbo Constitutional 92.22% 6.89 83.33% 100.00%
Default 100.00% 0.00 100.00% 100.00%
Default (flipped) 0.00% 0.00 0.00% 0.00%
PopAlign 100.00% 0.00 100.00% 100.00%
PairRM 100.00%
PairRM (tuned) 100.00%
Unaligned GPT-3.5 Turbo Constitutional 84.44% 9.81 73.33% 100.00%
Default 0.00% 0.00 0.00% 0.00%
Default (flipped) 100.00% 0.00 100.00% 100.00%
PopAlign 0.00% 0.00 0.00% 0.00%
PairRM 0.00%
PairRM (tuned) 100.00%
Table 5: Results for experiments on AlpacaEval data (65 samples). Averaged over 6 random seeds.
Dataset Model Annotator Mean Std Min Max
Aligned GPT-3.5-Turbo Constitutional 67.44% 3.43 63.08% 72.31%
Default 64.87% 1.16 63.08% 66.15%
Default (flipped) 33.08% 1.29 32.31% 35.38%
PopAlign 67.18% 0.79 66.15% 67.69%
GPT-4o Constitutional 68.46% 3.19 63.08% 72.31%
Default 72.56% 1.16 70.77% 73.85%
Default (flipped) 27.95% 1.80 26.15% 30.77%
PopAlign 69.05% 1.33 68.25% 71.43%
PairRM 64.62%
PairRM (tuned) 69.23%
Unaligned GPT-3.5-Turbo Constitutional 39.49% 3.32 35.38% 44.62%
Default 35.90% 1.26 33.85% 36.92%
Default (flipped) 66.67% 1.26 64.62% 67.69%
PopAlign 33.85% 2.57 30.77% 36.92%
GPT-4o Constitutional 66.41% 7.69 53.85% 72.31%
Default 26.92% 1.61 24.62% 29.23%
Default (flipped) 72.31% 1.69 70.77% 73.85%
PopAlign 30.24% 2.10 26.98% 32.26%
PairRM 35.38%
PairRM (tuned) 38.46%
Dataset Model Annotator Mean Std Min Max
Default 83.33% 0.00 83.33% 83.33%
User B constitution 83.33% 5.27 75.00% 91.67%
Default 88.89% 0.00 88.89% 88.89%
User B constitution 94.44% 6.09 88.89% 100.00%
Table 7: Results for cross-group experiments on PRISM data. Averaged over 6 random seeds.
Dataset Model Annotator Mean Std Min Max
Annotations Group A GPT-4o Group A constitution 77.22% 3.28 73.33% 83.33%
Default 58.33% 1.83 56.67% 60.00%
Group B constitution 50.00% 2.98 46.67% 53.33%
Annotations Group B GPT-4o Group A constitution 37.92% 2.04 35.00% 41.25%
Default 57.08% 1.02 56.25% 58.75%
Group B constitution 61.46% 4.50 55.00% 67.50%
Dataset Model Annotator Mean Std Min Max
Unaligned GPT-4o Default 26.15% 1.26 24.62% 27.69%
Constitutional 70.00% 0.89 69.23% 70.77%
Claude-3-Opus Default 24.23% 0.77 23.08% 24.62%
Constitutional 58.85% 0.77 58.46% 60.00%
Claude-3-Haiku Default 36.92% 0.00 36.92% 36.92%
Constitutional 59.65% 0.00 59.65% 59.65%
Dataset Model Annotator Mean Std Min Max
Constitutional (n=3) 75.00% 14.10 63.33% 100.00%
Constitutional (n=5) 86.67% 8.43 73.33% 96.67%
Constitutional (n=10) 90.00% 10.11 70.00% 96.67%
Table 10: Results for scaling experiments on unaligned AlpacaEval data (from 65 to 324 samples in test set). Averaged over 6 random seeds.
Dataset Model Annotator Mean Std Min Max
Original (65 samples) GPT-3.5-Turbo Default 35.90% 1.26 33.85% 36.92%
Constitutional 39.49% 3.32 35.38% 44.62%
GPT-4o Default 26.92% 1.61 24.62% 29.23%
Constitutional 66.41% 7.69 53.85% 72.31%
PairRM 37.35%
PairRM (tune) 46.60%
Large (324 samples) GPT-3.5-Turbo Default 45.83% 0.91 45.06% 46.91%
Constitutional 54.20% 1.56 52.01% 55.73%
GPT-4o Default 33.80% 0.58 33.02% 34.57%
Constitutional 61.47% 1.29 59.88% 62.96%
PairRM 37.35%
PairRM (tune) 50.00%
Appendix G Prompts
@@
Prompts are generally separated into two messages, a system message and a user message. We use the following format for all prompts (based on AlpacaEval’s formatting): <|im_start|> and <|im_end|> denote the start and end of a message, followed by the message type (system or user) and the content.
G.1 Principle generation
@@ G.3 Constitution evaluation
G.4 Non-constitutional baseline
We also evaluate the preferences the language model expresses when not given a constitution to follow, i.e., the biases inherent in the trained model when asked to select the “best” output. We adapted two of the default prompts from AlpacaEval for this purpose by removing references to an “instruction”, as this is not present in all pairwise comparison datasets. We selected the alpacaeval_gpt4_turbo_fn and chatgpt_fn prompts as they were reported to have the highest human agreement rate for the gpt-4-turbo and gpt-3.5-turbo models, respectively, while also being below an (estimated) price of 6\$/1k examples.  [^14] https://github.com/tatsu-lab/alpaca_eval/tree/v0.6.2/src/alpaca_eval/evaluators_configs
@@ G.4 Non-constitutional baseline
G.5 PopAlign baseline
The PopAlign baseline is based on the data generation approach by Wang et al. (2024) (described in more detail in Section F.8). To adapt the method for preference annotation, we combine both the bad and good generation prompt (taken from the elicitive contrast generation step in Table 7 by Wang et al. (2024)) into a single prompt. For a given response pair, this combined prompt asks to generate corresponding good and bad principles, and then asks to select the good response. We make this baseline available as a AlpacaEval annotator configuration as part of our package.
Listing 9: Original PopAlign prompt for generating good response, based on generated principles.
Please first consider the principles of crafting a good response, and then generate the response. Format your output as follows:
Thought: <Insights on creating a good response>
Response: <The good response>
Listing 10: Original PopAlign prompt for generating bad response, based on generated principles.
Please first consider the principles of crafting a bad response, and then generate the response. Format your output as follows:
Thought: <Insights on creating a bad response>
Response: <The bad response>
Listing 11: Our merged PopAlign preference annotation prompt.
<|im_start|>system
You are a helpful instruction-following assistant that selects responses.
<|im_end|>
<|im_start|>user
## Response A:
{output_1}
## Response B:
{output_2}
## Task
Please first consider the principles of crafting a good and a bad response, and then select the good response above. Format your output as follows:
Thought good: <Insights on creating a good response>
Thought bad: <Insights on creating a bad response>
Selected Response: <A or B>
## Your answer
<|im_end|>
Appendix H Constitutions
The following lists examples of full constitutions generated for each dataset and model combination in our experiments. To provide an unbiased view of the generated constitutions, we show the constitution with the highest, median, and lowest performance reconstruction accuracy on the experiment’s test set. Since even numbers of seeds are used, we chose the worse-performing constitution as a tie-breaker for the median.
H.1 Synthetic datasets
@@
H.1.1 Aligned
Listing 12: Best constitution on the ‘aligned’ synthetic dataset.
@@ H.1.1 Aligned
Listing 13: Median constitution on the ‘aligned’ synthetic dataset.
@@
Listing 14: Worst constitution on the ‘aligned’ synthetic dataset.
@@
H.1.2 Orthogonal
Listing 15: Best constitution on the ‘orthogonal’ synthetic dataset.
@@ H.1.2 Orthogonal
Listing 16: Median constitution on the ‘orthogonal’ synthetic dataset.
@@
Listing 17: Worst constitution on the ‘orthogonal’ synthetic dataset.
@@
H.1.3 Unaligned
Listing 18: Best constitution on the ‘unaligned’ synthetic dataset.
@@ H.1.3 Unaligned
Listing 19: Median constitution on the ‘unaligned’ synthetic dataset.
@@
Listing 20: Worst constitution on the ‘unaligned’ synthetic dataset.
@@
H.2 AlpacaEval datasets
H.2.1 Aligned
Listing 21: Best constitution on the ‘aligned’ AlpacaEval dataset.
@@ H.2.1 Aligned
Listing 22: Median constitution on the ‘aligned’ AlpacaEval dataset.
@@
Listing 23: Worst constitution on the ‘aligned’ AlpacaEval dataset.
@@
H.2.2 Unaligned
Listing 24: Best constitution on the ‘unaligned’ AlpacaEval dataset.
@@ 4. Select the response that lists key takeaways clearly and concisely
Listing 25: Median constitution on the ‘unaligned’ AlpacaEval dataset.
@@
Listing 26: Worst constitution on the ‘unaligned’ AlpacaEval dataset.
@@ H.3 Chatbot Arena
H.3.1 User A
Listing 27: Best constitution on User A annotations.
@@ H.3.1 User A
Listing 28: Median constitution on User A annotations.
@@
Listing 29: Worst constitution on User A annotations.
@@
H.3.2 User B
Listing 30: Best constitution on User B annotations.
@@ H.3.2 User B
Listing 31: Median constitution on User B annotations.
@@
Listing 32: Worst constitution on User B annotations.
@@
H.4 PRISM
Note that for personalized constitutions, we measure performance based on the ability to reconstruct the same group’s preferences. Due to the small number of samples, there is no split between test and training data.
H.4.1 Group A
Listing 33: Best constitution on Group A annotations.
1. Select the response that avoids redundancy and repetition.
2. Select the response that is concise and to the point.
3. Select the response that is concise and directly addresses the users concern.
4. Select the response that avoids unrelated information.
5. Select the response that provides a direct, concise answer.
Listing 34: Median constitution on Group A annotations.
1. Select the response that is concise and to the point.
2. Select the response that avoids unrelated information.
3. Select the response that avoids redundancy and repetition.
4. Select the response that is concise and directly addresses the users statement.
5. Select the response that provides a concise answer without offering additional details.
Listing 35: Worst constitution on Group A annotations.
1. Select the response that avoids irrelevant information.
2. Select the response that is concise and to the point.
3. Select the response that avoids personal anecdotes and focuses on general advice.
4. Select the response that avoids redundancy and repetition.
5. Select the response that provides a direct, concise answer.
H.4.2 Group B
Listing 36: Best constitution on Group B annotations.
1. Select the response that provides more detailed descriptions.
2. Select the response that avoids pretending to have human emotions.
3. Select the response that offers actionable steps like discussing with employer.
4. Select the response that asks for user preference on topics.
5. Select the response that mentions advanced technology and knowledge.
Listing 37: Median constitution on Group B annotations.
1. Select the response that provides a clear and factual explanation.
2. Select the response that provides more detailed steps.
3. Select the response that emphasizes proactive communication.
4. Select the response that emphasizes freshness and deliciousness.
5. Select the response that mentions the need for cross-checking information.
Listing 38: Worst constitution on Group B annotations.
1. Select the response that provides more actionable steps.
2. Select the response that mentions mindfulness and emotional awareness.
3. Select the response that provides a broader cultural context.
4. Select the response that mentions advanced technology and knowledge.
5. Select the response that emphasizes individual decision-making.
Appendix I Synthetic data generation